# FMNV: A Dataset of Media-Published News Videos for Fake News Detection

### v1

@@ Abstract
News media, particularly video-based platforms, have become deeply embedded in daily life, concurrently amplifying risks of misinformation dissemination. Consequently, multimodal fake news detection has garnered significant research attention. However, existing datasets predominantly comprise user-generated videos characterized by crude editing and limited public engagement, whereas professionally crafted fake news videos disseminated by media outlets—often politically or virally motivated—pose substantially greater societal harm. To address this gap, we construct FMNV, a novel dataset exclusively composed of news videos published by media organizations. Through empirical analysis of existing datasets and our curated collection, we categorize fake news videos into four distinct types. Building upon this taxonomy, we employ Large Language Models (LLMs) to automatically generate deceptive content by manipulating authentic media-published news videos. Furthermore, we propose FMNVD, a baseline model featuring a dual-stream architecture integrating CLIP and Faster R-CNN for video feature extraction, enhanced by co-attention mechanisms for feature refinement and multimodal aggregation. Comparative experiments demonstrate both the generalization capability of FMNV across multiple baselines and the superior detection efficacy of FMNVD. This work establishes critical benchmarks for detecting high-impact fake news in media ecosystems while advancing methodologies for cross-modal inconsistency analysis.
@@ 2 Related Work
3 Empirical Analysis
According to the existing communication studies [15, 16], the classification of fake news videos can be expanded from multiple dimensions, including production motivation, technical means, content nature and so on. We conduct an empirical analysis on the existing fake news video dataset and the fake news videos collected in the real world. Fake news often hides the truth by editing and splicing material from different sources to create misleading narratives, or by provoking emotional reactions through contradictory information. In view of the fact that news videos contain at least three modals of information: title, video clip and audio, we will analyze the semantic information differences among these three modes in detail. When the semantic information expressed by one of these modals is inconsistent with the others, or when the semantic information of the video is inconsistent with common sense, the video is more likely to be fake news. From the perspective of modal information, we divide fake news videos into the following four categories.
@@ 3 Empirical Analysis
Figure 1: An example of Contextual Dishonesty fake news video. The video shows children in Gaza longing for peace, but the title says Palestinians in Gaza anticipating continuation of hostilities.
3.1 Contextual Dishonesty
As the core information summary of the video, titles serve as a critical determinant of audience interpretation, yet deliberate fabrication of title modalities often creates semantic misalignment with actual content. Malicious title manipulation distorts narratives through selective omission or recontextualization of key contextual elements, employing strategies like decontextualized phrasing, inflammatory language, and sensationalized claims divorced from substantive information. For example, substituting the subject and object of an action, replacing neutral terminology with emotionally charged vocabulary that frames neutral events as conflicts, or omitting critical qualifying clauses that convert conditional statements into absolute declarations. Such tactics exploit audiences’ tendency to engage primarily with headlines, creating cognitive dissonance between the attention-grabbing title and the moderated substance within the content.
Figure 2: An example of Cherry-picked Editing fake news video. The video pieced together footage of Musk and General Motors, claiming that Musk was buying GM, but there was no visual evidence to prove it.
3.2 Cherry-picked Editing
As the core carrier of fake news, video clips can directly mislead the audience’s cognition through deliberate manipulation of visual elements. Malicious tampering of video clips primarily distorts meaning by selectively removing or omitting key scenes, dialogue, or context from the source material, and stitching together fragments of different events, places, or time periods to create a fake narrative. For example, by removing moderate remarks that make speakers appear extremist, by removing de-escalation scenes that portray peaceful protests as violent, or by piecing together pictures of different events to tie them together, such tactics exploit the audience’s inability to get the full context, amplifying bias and fueling misinformation.
Figure 3: An example of Synthetic Voiceover fake news video. The video only shows the prime minister at the party, but the audio is a voiceover overinterpretation that has nothing to do with the video itself.
3.3 Synthetic Voiceover
Synthetic voiceover fake news videos create misleading narratives by pairing AI-generated newscast speech with completely unrelated or repurposed video clips to create a fake illusion of authenticity. Using tools such as text-to-speech algorithms or voice-cloning software, creators overlay fake audio, such as fake political announcements, economic crises, or celebrity scandals, on unrelated visuals, such as unrelated clips of disasters or harmless public events. For example, an AI report claiming a "military conflict" might play over footage of a routine drill, or a fake health alert in an unrelated lab scenario. This mismatch exploits the audience’s tendency to believe in audiovisual coherence, masking the disconnect through authentic voice tone and professional editing, thus bypassing censorship.
3.4 Contrived Absurdity
Contrived absurdity fake news videos use seemingly consistent audio-visual content, artificial intelligent-generated video content or computer-generated special effects, and put forward illogical or absurd claims, such as "discovery of alien bodies" and "dog weighing 5 tons", relying on absurdity to attract the curiosity of the audience. The videos deliberately produce incredible scenarios, such as fake alien invasions or impossible scientific feats, with realistic editing to mimic legitimate news formats. Although on the surface several modes are consistent, the theme of their videos defies logic and common sense, aiming to attract viewers’ interest through absurd themes and thus gain attention.
@@ 3.4 Contrived Absurdity
4 The FMNV Dataset
In this chapter, we introduce the construction process of FMNV dataset in detail, including obtaining news videos published by the media from Twitter and YouTube, then using LLM for data augmentation according to the categories analyzed in Section 3. Finally, we perform data analysis on our dataset and compare it with other datasets.
@@ 5.1.3 Audio
6 Feature Aggregation
To better integrate the features of video frames and caption texts, we employ the cross-modal Transformer adopted by Qi et al. [3] to mutually enhance the information of each modalities. Specifically, our framework employs hierarchical co-attention for cross-modal fusion. First, title text features $F_{t}$ and audio-text features $F_{a}$ undergo co-attention:
$F_{t}^{\prime},F_{a}^{\prime}=\text{CoAtt}(F_{t},F_{a})$ (5)
The enhanced title features $F_{t}^{\prime}$ then interact with video frame features $F_{v}$ through:
@@ 6 Feature Aggregation
We concatenate $F_{t}^{\prime\prime}$, $F_{a}^{\prime}$, and $F_{v}^{\prime}$ into $F\in R^{3d}$, then project to $d$-dimensional space via learnable weights $W\in R^{d\times 3d}$. The fused features are processed by a Transformer encoder with $L$ layers:
$y=\text{Transformer}(FW^{\top})$ (7)
Finally, the fused features $y$ feeds into a classifier for prediction:
$\hat{y}=\text{MLP}(y)$ (8)
@@ 7.3 Experiments Results
7.4 Ablation Study
We further conduct a multimodal ablation study using FMNVD on the dataset to analyze the contributions of different modalities in FMNV. Specifically, we individually remove the title, video, and audio modalities from the data while simultaneously eliminating their corresponding cross-modal attention layers in the model. The experimental results presented in Table 4 demonstrate that all three modality ablations caused notable performance degradation in the baseline model. The relative importance of different modalities to model performance can be ranked as audio >title >frames. This phenomenon primarily stems from our implementation strategy where audio signals are first converted into text through Whisper for processing. The transcribed audio text inherently contains more linguistic content than video titles, and such extended textual information proves more impactful than shorter textual inputs (titles) and visual features in our framework.
Table 4: Results of ablation study.
@@ 7.4 Ablation Study
w/o Title w/o Frames w/o Audio 65.42 71.25 63.75 65.37 70.47 58.83 68.16 70.43 60.22 68.78 71.67 58.78

### v2

@@ Abstract
News media, particularly video-based platforms, have become deeply embedded in daily life, concurrently amplifying risks of misinformation dissemination. Consequently, multimodal fake news detection has garnered significant research attention. However, existing datasets predominantly comprise user-generated videos characterized by crude editing and limited public engagement, whereas professionally crafted fake news videos disseminated by media outlets—often politically or virally motivated—pose substantially greater societal harm. To address this gap, we construct FMNV, a novel dataset exclusively composed of news videos published by media organizations. Through empirical analysis of existing datasets and our curated collection, we categorize fake news videos into four distinct types. Building upon this taxonomy, we employ Large Language Models (LLMs) to automatically generate deceptive content by manipulating authentic media-published news videos. Furthermore, we propose FMNVD, a baseline model featuring a dual-stream architecture integrating CLIP and Faster R-CNN for video feature extraction, enhanced by co-attention mechanisms for feature refinement and multimodal aggregation. Comparative experiments demonstrate both the generalization capability of FMNV across multiple baselines and the superior detection efficacy of FMNVD. This work establishes critical benchmarks for detecting high-impact fake news in media ecosystems while advancing methodologies for cross-modal inconsistency analysis. Our dataset is available in https://github.com/DennisIW/FMNV.
@@ 2 Related Work
3 The FMNV Dataset
In this section, we construct a comprehensive dataset through a multi-stage process, integrating empirical analysis, real-world data collection, large-model-assisted augmentation, and in-depth dataset evaluation. First, we examine the semantic inconsistency patterns commonly found in fake news videos to guide data annotation. Then, we collect and curate video samples from various sources, followed by data augmentation using large language models and shot segmentation. Finally, we perform quantitative and qualitative analysis, comparing our dataset with existing benchmarks to highlight its quality and diversity.
3.1 Empirical Analysis
Existing communication studies [15, 16] suggest that fake news videos can be classified from multiple dimensions, such as production motives, technical means, and content nature. We analyze both public fake news video datasets and real-world collected samples. Fake news often misleads through edited and spliced footage or emotionally manipulative, contradictory content. Since news videos typically contain three modalities—title, video, and audio—we examine the semantic differences among them. When one modality conflicts with the others or defies common sense, the video is more likely to be fake. Based on modality-level inconsistency, we categorize fake news videos into four types:
@@ 3.1 Empirical Analysis
Figure 1: Example of Contextual Dishonesty: The video depicts children in Gaza longing for peace, while the title claims they anticipate continued conflict.
3.1.1 Contextual Dishonesty
Titles, as the video’s summary, significantly shape audience perception. Malicious title manipulation causes semantic dissonance by omitting context or using emotionally charged language. Common tactics include subject-object substitution, exaggeration, and removal of qualifiers—misleading readers who engage only with headlines.
Figure 2: Example of Cherry-picked Editing: A fabricated claim that Musk is buying GM, stitched together from unrelated clips.
3.1.2 Cherry-picked Editing
Video content is often altered by omitting key scenes or splicing unrelated footage to construct a false narrative. Tactics include removing moderate remarks, excluding peaceful scenes, or combining unrelated events, thereby amplifying bias and misleading viewers.
Figure 3: Example of Synthetic Voiceover: A party video overlaid with unrelated political commentary.
3.1.3 Synthetic Voiceover
These videos pair AI-generated or cloned speech with unrelated visuals. For example, a fake announcement might play over footage of a harmless event. The mismatch exploits viewers’ assumptions of audiovisual coherence, often bypassing detection through professional editing.
@@ 3.1.3 Synthetic Voiceover
3.1.4 Contrived Absurdity
Such videos use consistent multimodal content to present absurd claims (e.g., “alien bodies found” or “5-ton dog”), often generated by AI or CGI. Despite surface coherence, the narratives defy logic, aiming to attract attention through curiosity and shock.
@@ 4.1.3 Audio
Figure 6: The main structure of Co-Attention.
4.2 Feature Aggregation
To effectively integrate information across modalities, we adopt a hierarchical co-attention mechanism inspired by Qi et al. [3], the main structure of Co-Attention is shown in figure 6. which enables deep interaction between textual, auditory, and visual signals. This design helps the model capture fine-grained cross-modal alignments and contextual dependencies.
We first apply co-attention between the title features $F_{t}\in{R}^{L_{t}\times d}$ and the audio transcription features $F_{a}\in{R}^{L_{a}\times d}$. Both modalities are projected into a shared space using linear layers. Then, each modality attends to the other by computing multi-head attention in both directions:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (5)
Specifically, $F_{t}$ acts as query and attends to $F_{a}$ (as key and value), and vice versa. The outputs from both directions are integrated with residual connections and layer normalization:
$F_{t}^{\prime}$ $=\text{LayerNorm}(F_{t}+\text{Attn}_{a\rightarrow t})$ (6)
$F_{a}^{\prime}$ $=\text{LayerNorm}(F_{a}+\text{Attn}_{t\rightarrow a})$ (7)
After audio-text fusion, we use the enhanced title representation $F_{t}^{\prime}$ to interact with visual frame features $F_{v}\in{R}^{L_{v}\times d}$ using the same co-attention procedure. This produces updated features $F_{t}^{\prime\prime}$ and $F_{v}^{\prime}$:
@@ 4.2 Feature Aggregation
The final fused representation is obtained by concatenating the three enhanced modalities:
$F=\text{Concat}(F_{t}^{\prime\prime},F_{a}^{\prime},F_{v}^{\prime})\in{R}^{3d}$ (9)
This fused vector is linearly projected and passed into a Transformer encoder with $L$ layers to capture higher-order interactions, followed by a multi-layer perceptron (MLP) for final classification:
$\hat{y}=\text{MLP}(\text{Transformer}(FW^{\top}))$ (10)
By gradually enhancing the title features through audio and visual interactions, our model captures richer multi-modal semantics and improves its ability to detect subtle inconsistencies indicative of fake news.
@@ 5.3 Experiments Results
Table 4: Ablation study on FMNVD. (a) Modality removal. (b) Hyperparameter sensitivity.
(a) Modality Ablation
@@
w/o Title 65.42 65.37 68.16 68.78
w/o Frames 71.25 70.47 70.43 71.67
w/o Audio 63.75 58.83 60.22 58.78
(b) Hyperparameter Sensitivity
Setting Accuracy F1 Precision Recall
Patch size = 16 73.84 72.23 72.00 72.50
Seq. len. = 32 74.17 71.91 72.46 71.56
Seq. len. = 64 73.75 72.96 72.80 74.11
$\tau=0.3$ 74.17 71.12 72.82 70.44
$\tau=0.7$ 72.50 70.67 70.67 70.67
5.4 Ablation Study
To assess both the contribution of each modality and the robustness of our framework to key hyperparameters, we perform two sets of ablations on FMNVD:
(a) Modality Ablation We individually remove each modality (title, frames, audio) and its associated co-attention layers, and measure the resulting drop in performance. As shown in the top block of Table4, all three modalities are vital, with audio having the largest impact, followed by title and then visual frames. This phenomenon primarily stems from our implementation strategy where audio signals are first converted into text through Whisper for processing. The transcribed audio text inherently contains more linguistic content than video titles, and such extended textual information proves more impactful than shorter textual inputs (titles) and visual features in our framework.
(b) Hyperparameter Sensitivity To verify the stability of our detection framework, we further vary three core hyperparameters:
  • CLIP patch size (e.g. 16)
  • Maximum sequence length for text/audio features (e.g. 32 vs. 64 tokens)
  • Confidence threshold $\tau$ for filtering low-confidence predictions (e.g. 0.3 vs. 0.7)
The bottom block of Table 4 provides the performance under each setting, demonstrating that our model maintains stable accuracy and F1 across a reasonable range of each hyperparameter.